# Hypermodularity and community detection in higher-order networks

### v1

@@ II.3 Identity of the flattenings
To answer the question, start by explicitly writing the $k$-th standard flattening of $\mathbf{B}$:
@@
II.4 Building a vector equation
$\delta_{C_{i},C_{j}}=\frac{1}{2}\left(s_{i}s_{j}+1\right)\>,$ (9)
where $s_{i}$ are spin-like community-identifying variables that can take the values $+1$ and $-1$, depending on the community to which node $i$ is assigned. Next, note that, given $k$ nodes, the product of deltas in Eq. (3) does not need to include all $\binom{k}{2}$ possible combinations thereof. In fact, because of transitivity of equality, the product is best expressed as a chain, so that
Then, use Eq. (9) to explicitly rewrite the product chain for $k=3$:
@@ II.4 Building a vector equation
It becomes then clear that the vector expression for the hypermodularity of a bipartition, which we wrote in Eq. (17) for $k=3$, can be generalized to any $k$ as
where $\boldsymbol{\sigma}^{(k)}$ is a vector with $N^{k-1}$ components that can be written as
$\sigma_{\xi_{1},\xi_{2},\dotsc,\xi_{k-1}}=\sum_{\begin{subarray}{c}r=1\\ r\text{ odd}\end{subarray}}^{k-1}\left.\sum_{\begin{subarray}{c}\text{ordered $r$-choices }\left\{\varsigma_{1},\varsigma_{2},\dotsc,\varsigma_{r}\right\}\\ \text{from }\left\{\zeta_{1},\zeta_{2},\dotsc,\zeta_{k-1}\right\}\end{subarray}}\alpha_{1}\varsigma_{1}\varsigma_{2}\dotsb\varsigma_{r}\right|_{\zeta_{1}=s_{\xi_{1}},\zeta_{2}=s_{\xi_{2}},\dotsc,\zeta_{k-1}=s_{\xi_{k-1}}}\>,$ (23)
@@
II.5 Repeated bisections
When searching for the best partition of a network, it is clear that, in general, it may not consist of just two communities. A natural idea is then to apply the same method to the communities resulting from an initial partitioning, checking whether hypermodularity increases when splitting those even further. However, care must be taken when doing so. The reason is that a key step in the derivation of Eq. 21 and Eq. 23 relies on the vanishing of the sum of all elements of $\mathbf{B}$. However, bisecting an already existing community, this condition is no longer true, as one works with the subtensor corresponding to the nodes in the community, rather than with the data tensor of the entire network.
To find the correct way to deal with this situation, consider the change in hypermodularity that would result from a further split of a community $C$. This can be expressed as the sum of a positive term coming from the new split and a negative one, which is the contribution of the nodes in the community to the current value of the hypermodularity. From Eq. 3, it is clear that the former term consists of all and only the elements of the subtensor of $C$ that correspond to nodes assigned to the same community. Similarly, the latter term is the sum of all the element of the same subtensor. Thus, it is
@@ II.5 Repeated bisections
Now, write the product of deltas in terms of the spin-lile variables, as before. However, note that, for the reasons discussed above, the constant term can no longer be neglected. This results in
@@
which means that we can express the change in hypermodularity as the vector equation
@@
Note that the correction terms vanish if one considers the whole network, as expected. Thus, one can always directly use Eq. (28), since it reduces to Eq. (21) when the network has not yet been partitioned.
II.6 Refinement steps
After operating a bisection, we carry out a first refinement by computing how the hypermodularity would change if one were to switch the placement of each node from the community to which it was assigned to the other one. All the nodes are considered, and then, in a greedy fashion, the best move is accepted. The remaining nodes are then considered again, and the best move is accepted each time, until all the nodes have been evaluated. The best middle point throughout the procedure is then found. If it corresponds to an increase in hypermodularity, the moves up to that point are permanently accepted; otherwise, the state of the network is reverted to the one at the start of the procedure. The refinement is then repeated until no further improvement is obtained.
To compute the change in hypermodularity that each move would yield, first note that changing the assignment of a single node $i$ results in a single element of the vector $\mathbf{s}$ to switch sign. Similarly, the vector $\boldsymbol{\sigma}^{(k)}$ only changes in the elements that correspond to combinations of $k-1$ nodes involving node $i$. Thus, we can write $\boldsymbol{\sigma}^{(k)}\rightarrow\boldsymbol{\sigma}^{\prime(k)}=\boldsymbol{\sigma}^{(k)}+\boldsymbol{\Delta}\boldsymbol{\sigma}^{(k)}$ and $\mathbf{s}\rightarrow\mathbf{s}^{\prime}=\mathbf{s}+\boldsymbol{\Delta}\mathbf{s}$, where all the elements of $\boldsymbol{\Delta}\mathbf{s}$ are 0, except for the $i$-th one, which is equal to $-2s_{i}$. Then, write the hypermodularity of the partition after the change as
From this equation, it follows that
After all existing communities have been evaluated for further splits, and the local node-level refinement has been considered, we perform a global node-level optimization. The structure of this step is again that of an iterated greedy algorithm that evaluates the change in hypermodularity resulting from moving each node from the community to which it is currently assigned to any other possible community. As before, the best move is accepted, and the evaluation is repeated for the remaining nodes, until all nodes have been moved. Then, the middle point in the sequence of moves corresponding to the largest change in hypermodularity is found, and the all the moves up to that point are permanently accepted if such change is positive. The step is repeated until it produces an improvement in hypermodularity.
Finally, if the current partition consists of more than 1 community, a community-level optimization is carried out. As for the previous two steps, also this is an iterated greedy algorithm, repeated so long as the value of hypermodularity increases. The moves considered are the mergings of pairs of communities, akin to the main procedure of the Louvain algorithm [20]. Merging community $\alpha$ and community $\beta$ causes a change in hypermodularity that is due to all and only the combinations of $k$ nodes such that at least one is in community $\alpha$ and one is in community $\beta$. As in the previous step, each combination contributes an amount equal to the corresponding element of $\mathbf{B}$ divided by $k!m$. This means that the potential changes in hypermodularity can be represented as a matrix whose dimension is the current number of communities. When a move is accepted, it is convenient to take all the nodes belonging to the community with the higher index and assign them to the other one. This way, for all subsequent moves, the only matrix elements that change are those involving the community with the lower index, and, because of symmetry, only half a row and half a column need to be updated.
II.7 Numerical validation
To understand how this happens, consider a limit case of a hypergraph formed by $m$ isolated hyperedges. Its maximum hypermodularity, which is found by assigning all the nodes of each edge to a separate community, is $1-\frac{\left(k-1\right)!}{\left(km\right)^{k-1}}$, which, for large $m$, is approximately 1. However, the hypermodularity of the individual components, which consist each of a single edge, is identically 0. This shows that one must be very careful in using hypermodularity on networks that are not connected, and that in such cases the right approach is to treat each component as an independent network. Thus, our results also indicate that concerns of overfitting, as have been occasionally voiced also for traditional modularity, are mostly unfounded, and one can take the value of $0.2$ as a sort of physiological expected maximum hypermodularity of a random network, due to the non-vanishing probability that, even in a random structure, some sets of nodes may be more densely linked internally than they are between each other.
                Refer to caption Refer to caption                                                 Refer to caption
The primary school results, shown in Fig. 2, are quite revealing. When we use $k=2$, which corresponds to a traditional network, the algorithm divides the students into 5 communities, each corresponding to a different grade, with the exception that all the pupils of one of the classes of year 4 are assigned to the same community as the pupils of year 5. This partition, with a modularity of $0.3066$, is consistent with younger children tending to separate by age in their individual interactions, and with such boundaries becoming less rigid as they grow up. The communities of 3-people interactions ($k=3$) reveal a similar situation. Year-3 students keep to themselves, whereas one entire class of year-1 pupils joins the two of year-2 students to form a community. The pupils of the last two years are together in the same community. This division, with a hypermodularity of $0.71$, is again consistent with the dynamics just described. Year-3 students consider the two lower grades too young to form social interactions with, and they are themselves considered too young by the two higher grades. The data for $k=4$ show an intriguing division, with a hypermodularity of $0.7632$. Years 2 and 3 are joined in a single community, and years 1, 4 and 5 form another one, with two individual isolated students. A possible explanation for the former is that in looser groups of four people it is easier to include one or two members of different age without introducing too much disparity of maturity than it is in tighter groups of 2 or 3 people. The latter community is perhaps more striking. The age difference between the individuals in year 1 and those in year 4 and year 5 is of 3 to 4 years, and this community seems to result from the presence of siblings, who join the age-based social group of their brothers and sisters. Indeed, the average age gap between siblings in the period where the data were collected was, in fact, $3.5$ years [45]. Finally, for $k=5$ the network is almost completely fragmented, and the only nodes left with some edges separate into communities that are either uniform, or with at most one grade of difference between the members. Note that the network of 5-people interactions consists of 6 components, 5 of which contain a single hyperedge, and thus have a vanishing hypermodularity. The division of the only non-trivial component has a hypermodularity of $0.662$.
                            Refer to caption Refer to caption                Refer to caption
The analysis of the high-school data set, shown in Fig. 3, also provides insights into the social dynamics of the groups of students. When one considers only dyadic interactions, the network separates into communities that are completely identified not only by the topic studied, but also by the specific class in which the students are enrolled. This partition has a modularity of $0.5915$, and the only exception is that the mathematics and physics students form a single community, rather than splitting into their three classes. This evidently disproves the stereotype that sees mathematics and physics students as socially inept. In fact, in this study, they are the only group of people that socialize across classes, albeit in the same topic. When we take $k=3$, the network is very highly modular, with 3 large communities plus an individual node and a hypermodularity of $0.8313$. Here, the communities are entirely topic-based, but the students of psychology are joined with those of mathematics and physics. A possible explanation for this is that both physicists and psychologists study complex systems, even though using different means of representation. Thus, they have a larger affinity than one may imagine at a first thought. At the next larger size of edges, $k=4$, the meta-group of mathematics-physics-psychology students joins that of chemistry ones, with the notable exception of 9 mathematics-and-physics students who are now assigned to the community of biology. Also, the network starts fragmenting, with an isolated clique of 5 biology students. This partition of the large component has a hypermodularity of $0.8527$, and it is mostly caused by a few students who act as bridges between disciplines, and whose importance increases as the network has substantially fewer edges. For example, psychology student 325 is well connected with other psychology students and with mathematics-physics ones, chemistry student 284 is highly connected with a clique of psychology students and with numerous other chemistry ones, and mathematics-physics students 125 and 176 have several links to other students in the same class as well as to biology ones. Finally, when $k=5$ very few nodes remain with some edges. Five of them are a clique of chemistry students, and the other three components consist entirely of students of the same topic (two components for biology and one for chemistry). This suggests that these components are study groups, and the larger communities identified within them are their core members. It should be noted, nonetheless, that only the larger biology component is reasonably modular, with its partition corresponding to a hypermodularity of $0.3319$, whereas both other components have a hypermodularity of $0.1723$, which is below what one would expect from a random hypergraph, as discussed above.
III Conclusions
In this article, we have considered the task of detecting communities in networks with interactions beyond pairwise ones. Our way to attack this problem consists in defining a hypermodularity function for hypergraphs using a formalism that is conducive to the use of spectral methods to find the bisection of the network that maximizes it. Specifically, we have shown that this can be achieved by the use of higher-order singular value decomposition, if one rewrites the condition that the elements of a given set of nodes all belong to the same community as a separable expression of spin-like variables. To do so, we demonstrated a closed combinatorial expression that allows one to ultimately turn the hypermodularity equation into a vector form.
After implementing an algorithm based on these ideas, we applied it to random hypergraphs, finding a physiological value of maximum hypermodularity that one can expect to find in networks without an imposed community structure. This is due to the non-zero probability that, even placing all edges at random, one can end up with modules of nodes whose internal edges are denser than their external ones. Notably, the hypermodularity of random hypergraphs never exceeded $0.2$, suggesting that this value can be considered as a rough threshold to determine the presence of strong communities.
When applied to real-world data, our algorithm was able to detect communities at different levels of collective interaction. The communities thus identified can all be explained by assuming reasonable mechanisms for their formation. In fact, it can be argued that the communities found on social contact networks represent similarity of thought processes and reflect the biases of the individuals. Thus, we believe that our formulation of hypermodularity and our algorithm constitute powerful tools for the analysis of networks with higher-order interactions, and are highly valuable methods for extracting hidden information from data sets of diverse nature.
Acknowledgements.

### v2

@@ Appendix A Identity of the flattenings
To formally show why all the flattenings are identical, start by explicitly writing the $k$-th standard flattening of $\mathbf{B}$:
@@
Appendix B Building a vector equation
To turn Eq. (3) into a vector equation, first use Eq. (5) to explicitly rewrite the product chain for $k=3$:
@@ Appendix B Building a vector equation
It becomes then clear that the vector expression for the hypermodularity of a bipartition, which we wrote in Eq. (25) for $k=3$, can be generalized to any $k$ in the form given by Eqs. (7) and (9), which we rewrite here for the sake of completeness:
$Q=\frac{1}{2^{k-1}k!m}\mathbf{s}^{\mathrm{T}}\mathbf{E}\boldsymbol{\sigma}^{(k)}$ (29)
$\sigma^{(k)}_{\xi_{1},\xi_{2},\dotsc,\xi_{k-1}}=\sum_{\begin{subarray}{c}r=1\\ r\text{ odd}\end{subarray}}^{k-1}\left.\sum_{\begin{subarray}{c}\text{ordered $r$-choices }\left\{\varsigma_{1},\varsigma_{2},\dotsc,\varsigma_{r}\right\}\\ \text{from }\left\{\zeta_{1},\zeta_{2},\dotsc,\zeta_{k-1}\right\}\end{subarray}}\alpha_{1}\varsigma_{1}\varsigma_{2}\dotsb\varsigma_{r}\right|_{\zeta_{1}=s_{\xi_{1}},\zeta_{2}=s_{\xi_{2}},\dotsc,\zeta_{k-1}=s_{\xi_{k-1}}}\>,$ (30)
@@
Appendix C Repeated bisections
To find the expression for the correction to use when repeating the bisection process, start from Eq. (10), which we rewrite here for convenience:
@@ Appendix C Repeated bisections
Now, write the product of deltas in terms of the spin-like variables. However, note that, for the reasons discussed in Subsection II.2, the constant term can no longer be neglected. This results in
@@
Figure 4: Worked example of the hypermodularity maximization algorithm. (a) The network is a 12-node 3-uniform hypergraph consisting of three cliques, specifically on nodes 1–5, 5–8 and 9–12, joined by edges $(3,4,7)$, $(4,5,7)$, $(6,8,11)$ and $(8,9,11)$. All the edges within each clique are shown here as lines of different style. The edges that join the cliques are slightly thicker than the others. (b) After the first bisection, the best partition found consists of two communities, shown here in blue and pink, with a hypermodularity of $0.375$. (c) After the first local node-level refinement, the communities change, and the hypermodularity rises to $0.519531$. (d) Further passes of the method split the blue community into two, shown here in blue and emerald green. At this point, no further improvement is possible, and the algorithm stops. The hypermodularity of the final partition is $0.63623$.
@@
which means that we can express the change in hypermodularity as in Eqs. (11) and (12), which we rewrite below:
$\Delta Q=\frac{1}{2^{k-1}k!m}\mathbf{s}^{\mathrm{T}}\mathbf{E^{\prime}}\boldsymbol{\sigma}^{(k)}$ (35)
@@
Appendix D A worked example
Here, we provide an example of how our method works, by going step-by-step through the partitioning of a 12-node 3-uniform hypergraph. The hypergraph consists of three groups of nodes, namely nodes 1–4, 5–8 and 9–12, with all possible hyperedges of size 3 within each group. Additionally, as illustrated in Fig. 4 (a), edges $(3,4,7)$, $(4,5,7)$, $(6,8,11)$ and $(8,9,11)$ join the cliques.
In the beginning, all the nodes belong to the same community, and the hypermodularity is 0. Computing the left-singular vector corresponding to the largest singular value results in a partition where nodes 1, 2, 3, 4, 6 and 7 are assigned to community 1, and the remaining ones are assigned to community 2. Since this partition, depicted in Fig. 4 (b), has a hypermodularity of $0.375$, it is accepted.
Next, a local node-level refinement is attempted. The best set of moves consists in shifting nodes 5 and 8 from community 2 to community 1. These moves yield an increase in hypermodularity of $0.144531$. Thus, they are accepted, resulting in the partition shown in Fig. 4 (c), with hypermodularity of $0.519531$.
No global node-level refinement is possible, since it would be just a repetition of the last refinement step. Similarly, joining the two communities would make the hypermodularity vanish, and therefore no further modifications happen at this stage.
Trying to bisect community 1 using the higher-order SVD does not yield an increase in hypermodularity. Therefore, no bisection is performed on it. However, the best result from the local node-level refinement step consists in moving nodes 5–8 to a new community, which we label community 3. This move yields an increase of hypermodularity of $0.116699$. Thus, it is accepted, and the resulting partition, shown in Fig. 4 (d), has hypermodularity of $0.63623$.
Next, a bisection is attempted on community 2, but this would not result in an increase of hypermodularity, and therefore none is performed. Similarly, local and global node-level refinements and community-joining attempts all would decrease the value of hypermodularity, and no such moves are accepted. At this point, community 2 is marked as blocked, since no further improvement can be obtained by working solely on it.
Similar attempts at continuing the process, first on community 1 and then on community 3, do not produce any improvement. Thus, the whole procedure stops. The final result is that the method correctly split the original network into the three densely-connected modules, producing a partition with a significant value of hypermodularity.